OpenStudy (anonymous):

The function below shows the relationship between the length of each side of a cube (y) and the volume of the cube (x+3): y equals cube root of x plus 3

OpenStudy (anonymous):

@precal @SolomonZelman

OpenStudy (anonymous):

Hold on theres graphs

OpenStudy (anonymous):

@SolomonZelman @precal @jdoe0001 I dont understand

OpenStudy (solomonzelman):

Like \(\normalsize\color{blue}{ y=x^3+3 }\) or \(\normalsize\color{blue}{ y=(x+3)^3 }\) ??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

hold on ill show

OpenStudy (jdoe0001):

a quick screenshot of the material would be good

OpenStudy (solomonzelman):

oh yeah...

OpenStudy (anonymous):

\[y=^{3}\sqrt{x+3}\]

OpenStudy (anonymous):

The graph should cross the y axis between 1 and 2 because the cube root of 3 is in that range

OpenStudy (jdoe0001):

\(\large \bf y=\sqrt[3]{x+3}?\)

OpenStudy (anonymous):

yes @jdoe0001

OpenStudy (solomonzelman):

I think it is asking you for a function shift, Considering \(\normalsize\color{blue}{ y=\sqrt[3]{x} }\) When you say \(\normalsize\color{blue}{ y=\sqrt[3]{x\color{red}{ +v} } }\) you are shifting it v units left.

OpenStudy (anonymous):

and one more important tip would be checking where it passes a y value of 2. The corresponding x value is 5

OpenStudy (anonymous):

k

OpenStudy (solomonzelman):

graph \(\normalsize\color{blue}{ y=\sqrt[3]{x} }\) and shift it \(\normalsize\color{black}{ 3 }\) units left.

OpenStudy (anonymous):

choose a graph that meets the above criteria

OpenStudy (anonymous):

ok which one?

OpenStudy (anonymous):

the second one

OpenStudy (jdoe0001):

well... first off you'd need to graph or know what \(\Large y=\sqrt[3]{x}\) looks like

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

YAY

OpenStudy (jdoe0001):

well... so.. grab a few points.. say 5 points pick 5 random "x" to get the "y"

OpenStudy (anonymous):

THANK YOU @SolomonZelman @jdoe0001 @precal

OpenStudy (jdoe0001):

yw

OpenStudy (solomonzelman):

Yes. GOod that you want to learn, because otherwise you would have done this, https://www.desmos.com/calculator/psvtpyxt3o